# Towards Testing and Evaluating Vision-Language-Action Models for Robotic Manipulation: An Empirical Study

### v1

@@ 5.4. RQ4: Does the Change of Camera Pose Affect a VLA Model’s Performance?
Similar to RQ3, we re-executed the 1,434 passed test cases collected from RQ1 three times after randomly moving and rotating the camera from its default position ($N=4,302$). Specifically, we rotated the camera around each axis by an angle randomly sampled between $5^{\circ}$ and $5^{\circ}$. Additionally, we randomly moved the camera away from its center by a distance randomly sampled between 0 and 5 cm. We manually checked the corner cases where the camera was rotated by $5^{\circ}$ and moved by 5 cm and confirmed that the images captured with these camera angles still included the entire scene and objects. We report our experiment results in Table 4
@@
Overall, we found that the VLA model’s performance was greatly affected by the mutated camera poses. With mutated camera poses, the subject VLA models only passed 34.0% of the test cases that could be passed with default camera poses. These results indicate that the current VLA models are very sensitive to the camera’s extrinsic calibration results. When deploying VLA models, to achieve compatible performance, the developers need to carefully set up the camera poses, as in the robot demonstration data used for training. However, these may limit the generalizability of the VLA models. Future work may consider data augmentation to improve the VLA model’s robustness under different camera settings to enhance the model’s generalizability.
Current VLA models are not robust against mutated camera poses, resulting in degraded performance when the visual input is captured from an angle different from the default one. When rotating the camera for a maximum angle of 5 and moving it for a maximum distance of 5cm, the performance dropped to 34.0% of the performance on average with default camera poses.
Among seven subject VLA models, the most robust one, RT-1-400k, succeeded in 45.6% of the test cases with the mutated camera poses that were passed with the default camera poses. OpenVLA-7b also passed 31.3% of the test cases. We noticed a significant performance gap between two Octo models and other VLA models, where both two Octo models only passed less than 10% of the test cases while the other five models all passed more than 20% of the test cases. These results may have been largely attributed to the fact that Octo models used the smaller training dataset (~ 65K robot demonstrations) compared with the other ones (130K ~ 160K). As a result, the Octo model’s generalizability to visual inputs captured from different camera angles is significantly affected.
Among seven subject VLA models, RT-1-400k was the most robust one against the perturbation of camera poses. Octo-series models performed significantly less robustly than the other models. This may be largely attributed to the fact that Octo models were trained with only half of the robot demonstration data compared with the other subject models.
@@ 5.5. RQ5: How Robust Do VLA Models Perform Against Unseen Objects?
To investigate this research question, we used an external object dataset, YCB (Calli et al., 2015). YCB contains 56 objects that are not included in the Open-Embodiment-X dataset (the training/fine-tuning dataset of our subject VLA models). We leveraged VLATest to generate 1,000 testing scenes for each of the four subject robotic manipulation tasks while sampling the target object and the obstacles from the YCB dataset. We compared the performance of these subject VLA models with objects from YCB dataset to the results in RQ1.
Table 5. RQ5) Performance of subject VLA models with seen/unseen objects. Each cell in the column Seen and Unseen represents the success rate of different VLA models in different tasks. Darker the color, larger the negative differences between the performance on seen and unseen objects.
@@
Table 5 shows the performance of seven subject VLA models when manipulating seen/unseen objects. Overall, when manipulating with unseen objects, the performance of VLA models dropped significantly compared with manipulating with seen objects. The average performance across different VLA models dropped by 74.2%, 66.7%, 66.7%, and 20.0% in Task 1, Task 2, Task 3, and Task 4, respectively. Notably, we also found that none of these VLA models completely failed when manipulating unseen objects (except those that also failed with seen objects). These results suggest that, while all VLA models exhibit certain possibilities of generalizing themselves to unseen objects, the performance is by far unreliable.
@@
Current VLA models, though exhibit potential for generalization, are still unreliable for handling manipulation tasks with unseen objects. Our subject VLA models saw a performance drop ranging from 20.0% to 74.2% across four tasks when manipulating unseen objects compared with the seen objects.
@@
We also examined each testing scene and the success of the VLA model at every individual step, similar to the investigation of RQ1 and RQ2, to understand how unseen objects might have impacted its performance. Specifically, since the model must succeed at each step sequentially to complete the testing scene, we define a metric called the transfer rate ($Tr$), which calculates the model’s ability to transfer success from one step to the next, as follows.
(1) $\footnotesize Tr_{n}=\frac{Success\_rate_{n}}{Success\_rate_{n-1}},$
where $Success\_rate_{n}$ denotes the success rate of a task at step $n$. We define $Success\_rate_{0}=1000$.
@@
Figure 7. (RQ5) Transfer rate (TrTr) of each task when manipulating with seen/unseen objects. Entries with * mean that their mean differences are statistically significant ($p<0.05$).
@@
We present the results in Fig. 7. We also performed a paired $t$-test to examine the statistical differences between $Tr$ of seen and unseen objects. Overall, though $Tr_{1}$, $Tr_{2}$, and $Tr_{3}$ all decreased when manipulating with unseen objects instead of the seen ones, we only saw significant statistical differences on $Tr_{1}$ in Task 1 and Task 2 ($p=0.011$ and $p=0.007$, Cohen’s $d$ effect sizes are $1.34$ and $0.891$, respectively). These results reveal that the biggest challenge for current VLA models when dealing with unseen objects remains to accurately locate the correct object to manipulate, suggesting that the VLA models may not be able to recognize unseen objects in many cases. We did not observe significant statistical differences on $Tr_{1}$ in Task 3 and Task 4. A plausible explanation is due to the small number of passed test cases (~ 20 out of 1000) in both tasks.
@@ 6. Discussion
VLA models for robotic manipulation—too good to be true for now. While VLA models have the potential to revolutionize AI-enabled robotic manipulation, our experiment results in RQ1 and RQ2 reveal that the current models are still unreliable for common robotic manipulation tasks at the time of this study. As a result, deploying VLA models in high-stakes and safety-critical applications remains impractical.
@@
Our detailed analysis suggests that the deficiency mainly comes from the lack of capabilities in precisely interpreting complex task requirements and accurately localizing the correct target objects to manipulate. We believe there is a huge space for improvement in addressing these issues. One potential solution is to scale up the model size. Currently, even the largest model in our study, OpenVLA-7b, has only 7 billion parameters, which is significantly smaller than state-of-the-art (SOTA) models in other domains. For instance, one of the largest open-source LLMs, Llama 3.1, has 405 billion parameters. The closed-source LLMs such as GPT-4 and Claude-3.5 typically have even much higher numbers of model parameters. By scaling up the model, we may also observe emergent capabilities for VLA models (Brohan et al., 2023).
@@
In addition to scaling up the size of VLA models, exploring more effective prompting strategies could be another promising direction. Previous studies have demonstrated that improved prompting can significantly enhance the performance of LLMs across various tasks, such as solving mathematical problems (Wei et al., 2022; Yao et al., 2024) and code generation (Zhang et al., 2023; Chen et al., 2023). Our Finding 5.1 suggests that current VLA models struggle to decompose complex task instructions into multiple steps of action. To fix this, one may consider prompting the VLA model step by step. This also aligns with one of the popular prompting strategies in the field of LLMs – chain-of-thought prompting. While our study has only explored the most basic prompts, future research should explore more advanced prompting techniques to evaluate their impact on VLA model performance. Furthermore, one can also introduce multi-agent systems to split the robotic manipulation tasks among multiple VLA agents, which is a strategy widely used in other domains (Xia et al., 2024; Zhang et al., 2024; Yang et al., 2024).
@@
Addressing the robustness challenges. Our study results from RQ3 and RQ4 reveal that current VLA models lack robustness against several external factors, such as lighting conditions and camera poses. We also found that models with large-scale pre-training or those trained with larger robot demonstration datasets exhibit greater robustness compared to others. This highlights a promising research direction: enriching the robot demonstration data. Since manually collecting real-world robot demonstration data requires significant labeling efforts, researchers may leverage data augmentation techniques (Alomar et al., 2023; Bharadhwaj et al., 2024) or employing sim2real translation (Zhao et al., 2020; Höfer et al., 2021) to scale up training data by utilizing simulation environments.
@@
Moreover, our Finding 5.3, Finding 5.3, and Finding 5.4 show that when external factors like lighting and camera angles deviate significantly from default settings, VLA model performance drops accordingly. This suggests that existing robot demonstration datasets may lack diversity, especially regarding variations in external conditions. In future work, when collecting datasets for training/fine-tuning factors, practitioners should take these environmental factors into account, which may lead to the training of more robust VLA models.
@@
Assessing the capabilities of VLA models. Our Finding 5.5 and Finding 5.5 (RQ5) suggest that current VLA models struggled to perform tasks with unseen objects. In practice, it is unrealistic to expect VLA models to successfully perform manipulation for every possible task scenario. Thus, in parallel to the development of more powerful and robust VLA models, it is also important to design novel techniques to comprehensively assess the capabilities of VLA models and derive proper guidelines about the use of VLA models. To address these, potential solutions include throughout offline benchmarking and online risk assessment.
@@
In terms of offline benchmarking, our 18,604 generated test scenes across four tasks can serve as one of the early benchmarks for VLA models. In future work, practitioners may focus on expanding this benchmark to include a broader range of robotic manipulation tasks and diverse robot settings, covering various object types, environmental factors, and task complexities. In terms of online risk assessment, one may consider adopting SOTA techniques for risk assessment in other domains, such as uncertainty estimation (Huang et al., 2023b; Mukherjee and Awadallah, 2020) and safety monitoring (Xie et al., 2024). While general techniques may be directly used in the context of robotic manipulation with VLA models, it is unclear to what extent they can help with assessing the quality and reliability of decisions made by VLA models.
@@
Towards efficient testing for VLA models. Although our proposed VLATest successfully identified numerous failed test scenes across our subject VLA models, it also incurred significant time overheads. Specifically, we relied on a random sampler for sampling target objects and obstacles in the testing scenes, which may not have been the most efficient approach. Future research could focus on optimizing the $pose_{s}ampler$ in Algorithm 1 to strategically assign critical positions and orientations for objects and obstacles. For instance, one can consider metamorphic-based methods (Wang and Su, 2020; Yu et al., 2022; Gao et al., 2024) or search-based methods (Zhou et al., 2023; Cheng et al., 2023) to efficiently generate test scenes for VLA models. Meanwhile, researchers may also work on test prioritization (Li et al., 2024a; Dang et al., 2024; Feng et al., 2020) or test selection (Huang et al., 2024; Hu et al., 2023; Gao et al., 2022; Aghababaeyan et al., 2024) towards efficient testing for VLA models.

### v2

@@ 5.4. RQ4: Does the Change of Camera Pose Affect a VLA Model’s Performance?
Similar to RQ3, we re-executed the 1,434 passed test cases collected from RQ1 three times after randomly moving and rotating the camera from its default position ($N=4,302$). Specifically, we rotated the camera around each axis by an angle randomly sampled between $-5^{\circ}$ and $5^{\circ}$. Additionally, we randomly moved the camera away from its center by a distance randomly sampled between 0 and 5 cm. We manually checked the corner cases where the camera was rotated by $5^{\circ}$ and moved by 5 cm, and confirmed that the images captured with these camera angles still included the entire scene and objects. We report our experiment results in Table 5
@@
Overall, we found that the VLA model’s performance was greatly affected by the mutated camera poses. With mutated camera poses, the subject VLA models only passed 34.0% of the test cases that could be passed with default camera poses. These results indicate that the current VLA models are very sensitive to the camera’s extrinsic calibration results. When deploying VLA models, to achieve compatible performance, the developers need to carefully set up the camera poses, as in the robot demonstration data used for training. However, these may limit the generalizability of the VLA models. Future work may consider data augmentation to improve the VLA model’s robustness under different camera settings to enhance generalizability.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 8, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] Current VLA models are not robust against mutated camera poses, resulting in degraded performance when the visual input is captured from an angle that varies from the default one. When rotating the camera for a maximum angle of 5 and moving it for a maximum distance of 5cm, the performance dropped to 34.0% of the performance on average with default camera poses.
Among seven subject VLA models, the most robust one, RT-1-400k, succeeded in 45.6% of the test cases with the mutated camera poses that were passed with the default camera poses. OpenVLA-7b also passed 31.3% of the test cases. We noticed a significant performance gap between the two Octo models and other models, where both Octo models only passed less than 10% of the test cases, while the other five models all passed more than 20%. These results may have been largely attributed to the fact that Octo models used the smaller training dataset (~ 65K robot demonstrations) compared with the other ones (130K ~ 160K robot demonstrations). As a result, the Octo model’s generalizability to visual inputs captured from different camera angles is significantly affected.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 9, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] Among the seven subject VLA models, RT-1-400k was the most robust one against the perturbation of camera poses. Octo-series models performed significantly less robustly than the other models. This may be largely attributed to the fact that Octo models were trained with only half of the robot demonstration data compared with the other subject models.
@@ 5.5. RQ5: How Robust Do VLA Models Perform Against Unseen Objects?
To investigate this research question, we used an external object dataset, YCB (Calli et al., 2015). YCB contains 56 objects that are not included in the Open-Embodiment-X dataset (the training/fine-tuning dataset of our seven subject VLA models). Similar to RQ1, we leveraged VLATest to generate 1,000 testing scenes for each of the four subject robotic manipulation tasks while sampling the target object and the confounding objects from the YCB dataset. We compared the performance of these subject VLA models with objects from the YCB dataset to the results in RQ1.
Table 6. RQ5) Performance of subject VLA models with seen/unseen objects. Each cell in the column Seen and Unseen represents the success rate of different VLA models in different tasks. Darker the color, the larger the negative differences between the performance on seen and unseen objects.
@@
Table 6 shows the performance of seven subject VLA models when manipulating seen and unseen objects. Overall, when manipulating unseen objects, the performance of the subject VLA models dropped significantly compared with manipulating seen objects. The average performance across different VLA models dropped by 74.2%, 66.7%, 66.7%, and 20.0% in Task 1, Task 2, Task 3, and Task 4, respectively. Notably, we also found that none of these VLA models completely failed when manipulating unseen objects (except those that also failed with seen objects). These results suggest that, while all VLA models exhibit certain possibilities of generalizing themselves to unseen objects, the performance is by far unreliable.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 10, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ]
@@
Current VLA models, though exhibit potential for generalization, are still unreliable for handling unseen objects. Our subject VLA models saw a performance drop ranging from 20.0% to 74.2% across four tasks when manipulating unseen objects compared with the seen objects.
@@
We also examined each testing scene and the success of the VLA model at every individual step, similar to the investigation of RQ1 and RQ2, to understand how unseen objects might have impacted its performance. Specifically, since the model must succeed at each step sequentially to complete the testing scene, we define a metric called the transfer rate ($Tr$), which calculates the model’s ability to transfer success from one step to the next, as follows:
(1) $\small Tr_{n}=\frac{Success\_rate_{n}}{Success\_rate_{n-1}},$
where $Success\_rate_{n}$ denotes the success rate of a task at step $n$. We define $Success\_rate_{0}=100.0\%$.
@@
Figure 7. (RQ5) Transfer rate ($Tr$) of each task when manipulating with seen/unseen objects. Entries with * mean that their mean differences are statistically significant ($p<0.05$).
@@
We present the results in Fig. 7. We also performed a paired $t$-test to examine the statistical differences between $Tr$ of seen and unseen objects. Overall, though $Tr_{1}$, $Tr_{2}$, and $Tr_{3}$ all decreased when manipulating with unseen objects instead of the seen ones, we only saw significant statistical differences on $Tr_{1}$ in Task 1 and Task 2 ($p=0.011$ and $p=0.007$, Cohen’s $d$ effect sizes are $1.34$ and $0.891$, respectively). These results reveal that the biggest challenge for current VLA models when dealing with unseen objects remains to be accurately locating the correct object to manipulate, suggesting that the VLA models may not be able to recognize unseen objects in many cases. We did not observe significant statistical differences on $Tr_{1}$ in Task 3 and Task 4. A plausible explanation is due to the small number of passed test cases (~ 20 out of 1000) in both tasks.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 11, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ]
@@
5.6. RQ6: How Robust Do VLA Models Perform Against Task Instruction Mutations?
To study this research question, for each task, we prompted GPT-4o to generate 10 mutations of the standard task instruction (Sec. 4.4) that conveyed the same meanings. We only generated a limited number of mutations so that we could manually verify each mutation. Two of the authors manually validated and confirmed that all generated mutations were semantically equivalent to the corresponding standard task instruction. We then re-executed all test scenes generated in RQ1 (Sec. 5.1) for each of the four tasks. For each test scene, we randomly applied one of the mutated task instructions. These mutated instructions can be found in our Git repository.
Table 7. RQ6) Performance of subject VLA models with standard (Def.) and mutated (Mut.) task instructions. Each cell represents the success rate of different VLA models in different tasks. Darker the color, the larger the negative differences between the performance on standard and mutated task instructions.
VLA Models Task 1: Pick Up Task 2: Move Near Task 3: Put On Task 4: Put In
Def. Mut. Diff. Def. Mut. Diff. Def. Mut. Diff. Def. Mut. Diff.
RT-1-1k 0.7% 0.3% -57.1% 1.4% 0.9% -35.7% 0.0% 0.0% 0.0% 0.0%
RT-1-58k 28.2% 18.8% -33.3% 10.9% 6.4% -41.3% 0.0% 0.0% 0.0% 0.1%
RT-1-400k 34.4% 24.5% -28.7% 9.4% 8.8% -6.4% 0.5% 0.4% -20.0% 0.1% 0.1%
RT-1-X 19.5% 10.1% -48.2% 5.8% 5.6% -3.4% 2.3% 1.5% -34.8% 0.4% 0.5% +25.0%
Octo-small 0.8% 0.1% -87.5% 1.5% 1.5% 2.2% 2.1% -4.5% 1.1% 0.4% -63.6%
Octo-base 0.0% 0.3% 0.6% 0.4% -33.3% 1.2% 1.5% +25.0% 1.1% 0.5% -54.5%
OpenVLA-7b 5.9% 6.4% +8.5% 12.7% 12.2% -0.4% 2.1% 2.0% -4.8% 1.1% 2.0% +81.8%
Avg. 12.8% 8.6% -32.8% 6.0% 5.9% -1.7% 1.2% 1.1% -8.3% 0.5% 0.5%
  • $\dagger$ Diff. = (Mut. - Def.) / Def.
Table 7 presents the performance of seven subject VLA models when instructed with both the standard and mutated task instructions. When using the mutated task instructions, the average performance across the seven VLA models dropped by 32.8%, 1.7%, and 8.3% on Task 1, Task 2, and Task 3, respectively, while the performance difference on Task 4 was negligible.
We also found that larger models, such as OpenVLA-7b, were more robust against task instruction mutations. Specifically, OpenVLA-7b performed even better with the mutated task instructions on Task 1 and Task 4. For Task 2 and Task 3, the performance was only marginally affected. This robustness may be attributed to OpenVLA-7b’s use of Llama 2 as its language model, which potentially enhanced its language comprehension capabilities.
For instance, a mutated task instruction for Task 2 could be: place [OBJECT A] near [OBJECT B]. When prompted with this mutated instruction, OpenVLA-7b successfully manipulated 17.6% of the test scenes. In contrast, none of the other VLA models succeeded in more than 5% of the test scenes under the same instruction.
{mdframed}
[ nobreak=true, linecolor=black, roundcorner=12pt, backgroundcolor=gray!05, linewidth=0.3pt, leftmargin=0.5em, rightmargin=0.5em, topline=true, bottomline=true, frametitlerule=true, frametitlebackgroundcolor=gray!30, frametitlerulecolor=gray, frametitle=Finding 12, frametitleaboveskip=0.3em, frametitlebelowskip=0.35em, skipabove=5pt ] Current VLA models showed limited robustness against task instruction mutations. However, VLA models incorporating a large language model (e.g., OpenVLA-7b uses Llama 2-7b as the language model) are much more robust.
@@ 6. Discussion
VLA models for robotic manipulation—too good to be true for now. While VLA models have the potential to revolutionize robotic manipulation, our experiment results in RQ1 and RQ2 reveal that the current models are still unreliable for common robotic manipulation tasks at the time of this study. Deploying VLA models in high-stakes and safety-critical applications remains impractical.
@@
Our detailed analysis suggests that the deficiency mainly comes from the lack of capabilities in precisely interpreting complex task requirements and accurately localizing the correct target objects to manipulate. We believe there is a huge space for improvement in addressing these issues. One potential solution is to scale up the model size. Currently, even the largest model in our study, OpenVLA-7b, has only 7 billion parameters, which is significantly smaller than state-of-the-art (SOTA) models in other domains. For instance, one of the largest open-source LLMs, Llama 3.1, has 405 billion parameters. The closed-source LLMs such as GPT-4 and Claude-3.5 typically have even higher numbers of model parameters. By scaling up the model, we may also observe emergent capabilities for VLA models (Brohan et al., 2023).
@@
In addition to scaling up the size of VLA models, exploring more effective prompting strategies could be another promising direction. Previous studies have demonstrated that improved prompting can significantly enhance the performance of LLMs across various tasks, such as solving mathematical problems (Wei et al., 2022; Yao et al., 2023) and code generation (Zhang et al., 2023; Chen et al., 2024). Our Finding 5.1 suggests that current VLA models struggle to decompose complex task instructions into multiple steps of action. To fix this, one may consider prompting the VLA model step by step. This also aligns with one of the popular prompting strategies in the field of LLMs – chain-of-thought prompting. While our study has only explored the most basic prompts, future research should explore more advanced prompting techniques to evaluate their impact on VLA model performance.
Table 8. Different prompts for Task 4.
Model Original Prompt New Prompt
Grasp Move Success Grasp Move Success
RT-1-1k 1.1% 0.0% 0.0% 1.4% 0.0% 0.0%
RT-1-58k 1.3% 0.0% 0.0% 4.2% 0.6% 0.0%
RT-1-400k 8.9% 0.1% 0.1% 6.0% 0.1% 0.0%
RT-1-X 17.8% 0.7% 0.4% 28.0% 7.8% 0.4%
Octo-small 34.6% 1.1% 1.1% 30.0% 17.4% 0.0%
Octo-base 32.5% 1.1% 1.1% 22.1% 12.9% 0.2%
OpenVLA 19.9% 1.1% 1.1% 14.1% 4.4% 0.1%
To demonstrate the potential of improving VLA models’ performance through prompting, we crafted a new prompt for Task 4 by separating the task into two steps: pick up [object name] and then put [object name] into [object name]. We re-ran the 1000 test scenes generated in RQ1 (Sec. 5.1) for Task 4 using this new prompt and compared the performance with the original prompt. Our results in Table 8 show that with this new prompt, the success rate of grasping the correct object increased in three out of seven VLA models, while the success rate of moving the correct object improved in five out of seven models. The average success rate of moving the correct object increased from 0.6% to 6.2%. Although the overall task success rate did not increase, these results suggest that exploring prompting strategies for VLA models could be a promising research direction.
Furthermore, one can also introduce multi-agent systems to split the robotic manipulation tasks among multiple VLA agents, which is a strategy widely used in other domains (Xia et al., 2024; Zhang et al., 2024; Yang et al., 2024).
@@
Addressing the robustness challenges. Our study results from RQ3 and RQ4 reveal that current VLA models lack robustness against several external factors, such as lighting conditions and camera poses. We also found that models with large-scale pre-training or those trained with larger datasets of robot demonstration data exhibit greater robustness compared to others. This highlights a promising research direction: enriching the robot demonstration data. Since manually collecting real-world robot demonstration data requires significant labeling efforts, researchers may consider leveraging data augmentation techniques (Alomar et al., 2023; Bharadhwaj et al., 2024) or employing sim2real translation (Zhao et al., 2020; Höfer et al., 2021) to scale up training data by utilizing simulation environments. For instance, in the future, one could leverage a well-designed traditional controller to guide the robot in solving a test scene. The robot’s demonstration could then serve as data for re-training or fine-tuning.
@@
Moreover, our Finding 5.3, Finding 5.3, and Finding 5.4 show that when external factors like lighting and camera angles deviate significantly from default settings, VLA model performance drops accordingly. This suggests that existing robot demonstration datasets may lack diversity, especially regarding variations in external conditions. In future work, when collecting datasets for training/fine-tuning factors, practitioners should take these environmental factors into account, which may potentially lead to the training of more robust VLA models.
@@
Assessing the capabilities of VLA models. Our Finding 5.5 and Finding 5.5 (RQ5) suggest that current VLA models struggled to perform tasks with unseen objects. In practice, it is unrealistic to expect VLA models to successfully perform manipulation for every possible task scenario. Thus, in parallel to the development of more powerful and robust VLA models, it is also important to design novel techniques to comprehensively assess the capabilities of VLA models and derive proper guidelines about the use of VLA models. To address these, potential solutions include offline benchmarking and online risk assessment.
@@
In terms of offline benchmarking, our 18,604 generated test scenes across four tasks can serve as one of the early benchmarks for VLA models. In future work, practitioners may focus on expanding this benchmark to include a broader range of robotic manipulation tasks and diverse robot settings, covering various object types, environmental factors, and task complexities. In terms of online risk assessment, one may consider adopting SOTA techniques for risk assessment in other domains, such as uncertainty estimation (Huang et al., 2025; Mukherjee and Awadallah, 2020) and safety monitoring (Xie et al., 2024). While general techniques may be directly used in the context of robotic manipulation with VLA models, it is unclear to what extent they can help with assessing the quality and reliability of decisions made by VLA models.
@@
Towards efficient testing for VLA models. Although our proposed VLATest successfully identified numerous failed test scenes across our subject VLA models, it also incurred significant time overheads. Specifically, we relied on a random sampler for sampling target objects and confounding objects in the testing scenes, which may not have been the most efficient approach. Future research could focus on optimizing the $pose_{s}ampler$ in Algorithm 1 to strategically assign critical positions and orientations for objects and confounding objects. For instance, one can consider metamorphic-based methods (Wang and Su, 2020; Yu et al., 2022; Gao et al., 2024) or search-based methods (Zhou et al., 2023; Cheng et al., 2023) to efficiently generate test scenes for VLA models. Meanwhile, researchers may also work on test prioritization (Li et al., 2024a; Dang et al., 2024; Feng et al., 2020) or test selection (Huang et al., 2024; Hu et al., 2023; Gao et al., 2022; Aghababaeyan et al., 2024) towards efficient testing for VLA models.